half of henry's age added to one-third of daisy's age is 11 years. Six years from now, the sum of their ages will be 40 years. how old is each?
Answer by Jk22(389)

half of henry's age added to one-third of daisy's age is 11 years. Six years from now, the sum of their ages will be 40 years. how old is each?
let h=henry's age, d=daisi's
1) h/2+d/3=11
2) (h+6)+(d+6)=40 => h+d=28 => h=28-d
->1 : 14-d/2+d/3=14-d/6=11 => d/6=3 => d=18 years=daisi's age
henry's age h=28-18=10 years